Grade 12th passZoology

Equal weight of CH4 and H2 is taken in a vessel. The fraction of total pressure exerted by H2 Molecules is
A. 1/2
B. 1/8
C. 8/9
D. 1/9

Pushpa Raj Joshi
6 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
Hiii 
Let the equal mass to be taken is 16 g . 
Then no. Of moles of CH4 = 16/16 = 1
No. Of moles of H2 = 16/2 =8 ( no.of moles =given weight/molecular weight)
Then mole fraction of H2 = 8/(8+1)
= 8/9 
Partial pressure caused by H2 = 8/9(total pressure)
As it is done in container so total pressure = 1 ATM
Partial pressure caused by H2 = 8/9×1
=8/9